OpenStudy (thedj4jc):

The center of a circle is at the origin. An endpoint of a diameter of the circle is at (-3, -4). How long is the diameter of the circle? 5 10 25

OpenStudy (anonymous):

What's the distance between the given point and the origin?

OpenStudy (anonymous):

(0,0) is the center , thats simple count the distance between the two points , \[distance = \sqrt{(-3 - 0 )² + ( -4 - 0 ) ²}\] \[distance = \sqrt{9 + 16 }\] \[distance = \sqrt{25 }\] \[distance = 5\]

OpenStudy (anonymous):

\[\sqrt{(x2-x1)² + (y2- y1)²}\]

OpenStudy (thedj4jc):

Thank you!!!!!!!